Bobs Bungee Jumping Inc. has estimated that ½ of their 1,000 customers will be injured in the coming year. The expected severity: occurrence is $1,000. What is the expected loss per exposure? O $10,000 $500 O $10 O $100 O $50,000
Answer #1

Correct answer is option 500

Number of customer 1000
Estimation of injury = 1/2*1000 500
Expected cost = 1000
Total exposure =500*1000 500000
Expected loss per exposure =500000/1000 500